# Twisted character formula for toral supercuspidal representations

### v1

@@ Abstract.
We establish an explicit formula for twisted Harish-Chandra characters of toral supercuspidal representations of $p$-adic reductive groups under several technical assumptions. Our setup especially includes the case of a quasi-split group equipped with a involution.
@@ 1. Introduction
Now let us state our main result. The setup is as follows. Let $\mathbf{G}$ be a tamely ramified quasi-split connected reductive group over $F$ and $\theta$ an $F$-rational involution of $\mathbf{G}$ preserving an $F$-splitting of $\mathbf{G}$. We assume that the restricted root system with respect to $(\mathbf{G},\mathbf{S},\theta)$ does not contain any restricted root of type $2$ or $3$. We also assume a technical condition on the restricted root system, which is satisfied when $\theta$ is a involution, for example (see Section 6.4). Let $\tilde{\mathbf{S}}\subset\tilde{\mathbf{G}}$ be a tame elliptic twisted maximal torus (see Section 3.2 for the definition of a twisted maximal torus); by fixing a base point $\underline{\eta}\in\tilde{\mathbf{S}}(F)$, we may write $\tilde{\mathbf{S}}=\mathbf{S}\cdot\underline{\eta}$ for a tame elliptic maximal torus $\mathbf{S}$ of $\mathbf{G}$. The $\underline{\eta}$-conjugation on $\mathbf{G}$ preserves $\mathbf{S}$, hence induces an $F$-rational automorphism $\theta_{\mathbf{S}}$ of $\mathbf{S}$. We consider a toral pair $(\mathbf{S},\vartheta)$ such that $\vartheta=\vartheta\circ\theta_{\mathbf{S}}$. Then the corresponding toral supercuspidal representation $\pi:=\pi_{(\mathbf{S},\vartheta)}$ is a $\theta$-stable toral supercuspidal representation. Let $r$ be the depth of $\pi$.
@@
$V=\bigoplus_{\alpha\in\dot{\Phi}}V_{\alpha},$
@@
where the action of $S$ on $V_{\alpha}$ is given by $\alpha$ (each $V_{\alpha}$ could be zero). We call $\alpha\in\Phi$ symmetric (resp. asymmetric) if $\alpha\in\Gamma\alpha$ (resp. $\alpha\notin\Gamma\alpha$). When $\alpha$ is symmetric, $V_{\alpha}$ is a symplectic subspace. When $\alpha$ is asymmetric, $V_{\alpha}\oplus V_{-\alpha}$ forms a symplectic subspace. Furthermore, the resulting decomposition
$V=\bigoplus_{\Sigma\alpha\in\ddot{\Phi}}V_{\Sigma\alpha}$
@@
is orthogonal, where $\Sigma:=\Gamma\times\{\pm 1\}$, $\ddot{\Phi}:=\Phi/\Sigma$, and $V_{\Sigma\alpha}=V_{\alpha}$ or $V_{\Sigma\alpha}=V_{\alpha}\oplus V_{-\alpha}$ according to whether $\alpha$ is symmetric or asymmetric. This implies that, noting that the conjugate action of $s$ on $V$ preserves each $V_{\Sigma\alpha}$, we can compute $\operatorname{tr}(\omega([s]))$ as the product of the characters of Heisenberg–Weil representations $\omega_{\Sigma\alpha}$ of $H(V_{\Sigma\alpha})$ for $\Sigma\alpha\in\ddot{\Phi}$:
$\operatorname{tr}(\omega([s]))=\prod_{\Sigma\alpha\in\ddot{\Phi}}\operatorname{tr}(\omega_{\Sigma\alpha}([s])).$
@@
Each symplectic space $V_{\Sigma\alpha}$ is enough understandable, so that the character $\operatorname{tr}(\omega_{\Sigma\alpha}([s]))$ can be explicated using [Gér77]. For example, when $\alpha$ is asymmetric, $V_{\alpha}$ and $V_{-\alpha}$ form a polarization of $V_{\Sigma\alpha}$ and $[s]$ preserves this polarization. In fact, Gérardin’s character formula for such symplectic automorphism becomes particularly simple. Consequently, we obtain a certain sign function constituting $\varepsilon$ of the ADS formula (Theorem 1.1).
@@
Now let us go back to the twisted situation, that is, the computation of $\Theta_{\tilde{\rho}}(\eta)$ for $\eta\in\tilde{S}:=\tilde{\mathbf{S}}(F)$. Again in this situation, the conjugate action of $\eta$ on $J$ preserves the symplectic structure of $V$, hence it is enough to compute $\operatorname{tr}(\omega([\eta]))$ for the symplectic automorphism $\eta]$ of $V$ induced by $\eta$. The difference from the untwisted case is that $\eta]$ does not necessarily acts on each symplectic subspace $V_{\Sigma\alpha}$. To be more precise, note that the action of $\eta$ on $V$ induces a permutation on the set of roots $\Phi$. If we write $\theta\colon\Phi\rightarrow\Phi$ for this permutation by abuse of notation (this does not depend on $\tilde{\eta}\in\tilde{S}$), then $\eta]$ maps $V_{\Sigma\alpha}$ onto $V_{\Sigma\theta(\alpha)}$. Therefore, the smallest pieces of an orthogonal decomposition of $V$ we can utilize in the twisted setting are not individual $V_{\Sigma\alpha}$’s, but their sums $V_{\Sigma\alpha}+V_{\Sigma\theta(\alpha)}+V_{\Sigma\theta^{2}(\alpha)}+\cdots$. Accordingly, the product expression of $\operatorname{tr}(\omega([\eta]))$ is of the form
$\operatorname{tr}(\omega([\eta]))=\prod_{\alpha\in\Theta\backslash\ddot{\Phi}}\operatorname{tr}(\omega_{\Sigma\alpha}([\eta]^{m_{\alpha}})),$
@@
In fact, this is exactly the source of the complication. For example, let us consider the case where $\alpha$ is asymmetric (hence $V_{\Sigma\alpha}=V_{\alpha}\oplus V_{-\alpha}$ gives a polarization) and moreover $V_{\Sigma\theta(\alpha)}=V_{\Sigma\alpha}$. In this case, we have $\theta(\Gamma\alpha)=\pm\Gamma\alpha$. If $\theta(\Gamma\alpha)=\Gamma\alpha$, then $V_{\alpha}$ and $V_{-\alpha}$ are preserved by $\eta]$, thus the character of $\eta]$ can be computed easily as in the untwisted case. However, it could also happen that $\theta(\Gamma\alpha)=-\Gamma\alpha$, then $V_{\alpha}$ and $V_{-\alpha}$ are swapped by $\eta]$. This means that the symplectic automorphism $\eta]$ of $V_{\Sigma\alpha}$ no longer preserves the polarization, which makes the description of Gérardin’s formula complicated.
What is happening in this example is that the $\theta$-orbit of $\alpha$ is symmetric with respect to the $\Gamma$-action even though $\alpha$ itself is asymmetric. Therefore, to compute each $\operatorname{tr}(\omega_{\Sigma\alpha}([\eta]^{m_{\alpha}}))$, we have to care about the relationship between the symmetry of the root system $\Phi$ and the twist $\theta$. What we give eventually in this paper is to present an “algorithm” to determine each $\operatorname{tr}(\omega_{\Sigma\alpha}([\eta]^{m_{\alpha}}))$ based on a case-by-case argument on the type of $\alpha$. Hence, we do not really give a complete description of the twisted character in general. However, depending on the situation ($\mathbf{G}$, $\theta$, and $\mathbf{S}$), the algorithm may terminate rather quickly. For example, it is the case when $\theta$ is involutive.
@@ 2.1.1. $p$-adic fields
Let $p$ be a prime number. We fix a $p$-adic field $F$, i.e., $F$ is a finite extension of $\mathbb{Q}_{p}$. We fix an algebraic closure $\overline{F}$ of $F$ and always work within $\overline{F}$ when we consider algebraic extensions of $F$. For any finite extension $E$ of $F$, we write $\mathcal{O}_{E}$, $\mathfrak{p}_{E}$, $k_{E}$, and $\Gamma_{E}$ for the ring of integers of $E$, the maximal ideal of $\mathcal{O}_{E}$, the residue field, and the absolute Galois group $\operatorname{Gal}(\overline{F}/E)$ of $E$, respectively. When $E=F$, we omit the subscript $F$ from these symbols. For any finite extension $E$ of $F$, we write $W_{E}$, $I_{E}$, and $P_{E}$ for the Weil groups of $E$, its inertia subgroup, and its wild inertia subgroup, respectively. For any $r\in\mathbb{R}_{>0}$, let $I_{F}^{r}$ denote the $r$-th upper ramification filtration of $I_{F}$. We fix a valuation $\operatorname{val}_{F}$ of $F$ such that $\operatorname{val}_{F}(F^{\times})=\mathbb{Z}$. We extend it to $\overline{F}$ and again write $\operatorname{val}_{F}$ for it. We define an absolute value $|\cdot|_{\overline{F}}$ of $\overline{F}$ by $|\cdot|_{\overline{F}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}q^{-\operatorname{val}_{F}(\cdot)}$, where $q$ is the cardinality of the residue field $k_{F}$.
@@ 2.1.6. Finite sets with Galois actions
We put $\Sigma\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\Gamma\times\{\pm 1\}$. Suppose that $\Phi$ is a finite set with an action of $\Sigma$, e.g., the set of roots of an $F$-rational maximal torus in a connected reductive group (1-1 acts on $\Phi$ via $\alpha\mapsto-\alpha$ in this case). Following [AS09], we put $\dot{\Phi}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\Phi/\Gamma$ and $\ddot{\Phi}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\Phi/\Sigma$.
@@
Note that, if $\alpha$ is symmetric, $\Gamma\alpha=\Sigma\alpha$. This implies that the sets $\dot{\Phi}_{\mathrm{sym}}$ and $\ddot{\Phi}_{\mathrm{sym}}$ can be naturally identified (and, of course, the same is true for $\Phi_{\mathrm{ur}}$ or $\Phi_{\mathrm{ram}}$).
@@ 3.2. Twisted maximal torus
    for any $\eta\in\tilde{\mathbf{S}}$, $\mathbf{S}^{\natural}$ is a maximal torus of $\mathbf{G}_{\eta}$.
@@ 4.1. Regular supercuspidal representations
The “fibers” of Yu’s construction was investigated by Hakim–Murnaghan; in [HM08], they introduced an equivalence relation called $\mathbf{G}$-equivalence and proved that two cuspidal $\mathbf{G}$-data give rise to isomorphic supercuspidal representations if and only if two data are $\mathbf{G}$-equivalent.
@@ 4.2. Toral supercuspidal representations
Since the depth of $\vartheta$ is $r$, we can extend $\vartheta$ to a character $\hat{\vartheta}$ of $G_{\mathbf{x},s+}$ satisfying $\hat{\vartheta}|_{(S,G)_{\mathbf{x},(r+,s+)}}\equiv\mathbbm{1}$. Then, by the definition of the $\mathbf{G}$-genericity, there exists an element $X^{\ast}$ of $\mathfrak{s}^{\ast}_{-r}$ which is $\mathbf{G}$-generic of depth $r$ in the sense of [Yu01, Section 8] and satisfies
@@ 5.1. Twisted character of a $\theta$-stable representation
$\tilde{\pi}(f)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\int_{\delta\in\tilde{G}}f(\delta)\tilde{\pi}(\delta)\,d\delta,$

### v2

@@ Abstract.
We establish an explicit formula for twisted Harish-Chandra characters of toral supercuspidal representations of $p$-adic reductive groups under several technical assumptions. Our setup especially includes the case of a quasi-split group equipped with an involution.
@@ 1. Introduction
Now let us state our main result. The setup is as follows. Let $\mathbf{G}$ be a tamely ramified quasi-split connected reductive group over $F$ and $\theta$ an $F$-rational involution of $\mathbf{G}$ preserving an $F$-splitting of $\mathbf{G}$. We assume that the restricted root system with respect to $(\mathbf{G},\mathbf{S},\theta)$ does not contain any restricted root of type $2$ or $3$. We also assume a technical condition on the restricted root system, which is satisfied when $\theta$ is an involution, for example (see Section 6.4). Let $\tilde{\mathbf{S}}\subset\tilde{\mathbf{G}}$ be a tame elliptic twisted maximal torus (see Section 3.2 for the definition of a twisted maximal torus); by fixing a base point $\underline{\eta}\in\tilde{\mathbf{S}}(F)$, we may write $\tilde{\mathbf{S}}=\mathbf{S}\cdot\underline{\eta}$ for a tame elliptic maximal torus $\mathbf{S}$ of $\mathbf{G}$. The $\underline{\eta}$-conjugation on $\mathbf{G}$ preserves $\mathbf{S}$, hence induces an $F$-rational automorphism $\theta_{\mathbf{S}}$ of $\mathbf{S}$. We consider a toral pair $(\mathbf{S},\vartheta)$ such that $\vartheta=\vartheta\circ\theta_{\mathbf{S}}$. Then the corresponding toral supercuspidal representation $\pi:=\pi_{(\mathbf{S},\vartheta)}$ is a $\theta$-stable toral supercuspidal representation. Let $r$ be the depth of $\pi$.
@@
$V=\bigoplus_{\dot{\alpha}\in\dot{\Phi}}V_{\alpha},$
@@
where the action of $S$ on $V_{\alpha}$ is given by $\alpha$ (each $V_{\alpha}$ could be zero). We call $\alpha\in\Phi$ symmetric (resp. asymmetric) if $\alpha\in\dot{\alpha}=\Gamma\alpha$ (resp. $\alpha\notin\dot{\alpha}=\Gamma\alpha$). When $\alpha$ is symmetric, $V_{\alpha}$ is a symplectic subspace. When $\alpha$ is asymmetric, $V_{\alpha}\oplus V_{-\alpha}$ forms a symplectic subspace. Furthermore, the resulting decomposition
$V=\bigoplus_{\ddot{\alpha}\in\ddot{\Phi}}V_{\ddot{\alpha}}$
@@
is orthogonal, where $\Sigma:=\Gamma\times\{\pm 1\}$, $\ddot{\Phi}:=\Phi/\Sigma$, and $V_{\ddot{\alpha}}=V_{\alpha}$ or $V_{\ddot{\alpha}}=V_{\alpha}\oplus V_{-\alpha}$ according to whether $\alpha$ is symmetric or asymmetric. This implies that, noting that the conjugate action of $s$ on $V$ preserves each $V_{\ddot{\alpha}}$, we can compute $\operatorname{tr}(\omega([s]))$ as the product of the characters of Heisenberg–Weil representations $\omega_{\ddot{\alpha}}$ of $H(V_{\ddot{\alpha}})$ for $\ddot{\alpha}\in\ddot{\Phi}$:
$\operatorname{tr}(\omega([s]))=\prod_{\ddot{\alpha}\in\ddot{\Phi}}\operatorname{tr}(\omega_{\ddot{\alpha}}([s])).$
@@
Each symplectic space $V_{\ddot{\alpha}}$ is enough understandable, so that the character $\operatorname{tr}(\omega_{\ddot{\alpha}}([s]))$ can be explicated using [Gér77]. For example, when $\alpha$ is asymmetric, $V_{\alpha}$ and $V_{-\alpha}$ form a polarization of $V_{\ddot{\alpha}}$ and $[s]$ preserves this polarization. In fact, Gérardin’s character formula for such symplectic automorphism becomes particularly simple. Consequently, we obtain a certain sign function constituting $\varepsilon$ of the ADS formula (Theorem 1.1).
@@
Now let us go back to the twisted situation, that is, the computation of $\Theta_{\tilde{\rho}}(\eta)$ for $\eta\in\tilde{S}:=\tilde{\mathbf{S}}(F)$. Again in this situation, the conjugate action of $\eta$ on $J$ preserves the symplectic structure of $V$, hence it is enough to compute $\operatorname{tr}(\omega([\eta]))$ for the symplectic automorphism $\eta]$ of $V$ induced by $\eta$. The difference from the untwisted case is that $\eta]$ does not necessarily acts on each symplectic subspace $V_{\ddot{\alpha}}$. To be more precise, note that the action of $\eta$ on $V$ induces a permutation on the set of roots $\Phi$. If we write $\theta\colon\Phi\rightarrow\Phi$ for this permutation by abuse of notation (this does not depend on $\tilde{\eta}\in\tilde{S}$), then $\eta]$ maps $V_{\ddot{\alpha}}$ onto $V_{\Sigma\theta(\alpha)}$. Therefore, the smallest pieces of an orthogonal decomposition of $V$ we can utilize in the twisted setting are not individual $V_{\ddot{\alpha}}$’s, but their sums $V_{\ddot{\alpha}}+V_{\Sigma\theta(\alpha)}+V_{\Sigma\theta^{2}(\alpha)}+\cdots$. Accordingly, the product expression of $\operatorname{tr}(\omega([\eta]))$ is of the form
$\operatorname{tr}(\omega([\eta]))=\prod_{\ddot{\alpha}\in\Theta\backslash\ddot{\Phi}}\operatorname{tr}(\omega_{\ddot{\alpha}}([\eta]^{m_{\alpha}})),$
@@
In fact, this is exactly the source of the complication. For example, let us consider the case where $\alpha$ is asymmetric (hence $V_{\ddot{\alpha}}=V_{\alpha}\oplus V_{-\alpha}$ gives a polarization) and moreover $V_{\Sigma\theta(\alpha)}=V_{\ddot{\alpha}}$. In this case, we have $\theta(\dot{\alpha})=\pm\dot{\alpha}$. If $\theta(\dot{\alpha})=\dot{\alpha}$, then $V_{\alpha}$ and $V_{-\alpha}$ are preserved by $\eta]$, thus the character of $\eta]$ can be computed easily as in the untwisted case. However, it could also happen that $\theta(\dot{\alpha})=-\dot{\alpha}$, then $V_{\alpha}$ and $V_{-\alpha}$ are swapped by $\eta]$. This means that the symplectic automorphism $\eta]$ of $V_{\ddot{\alpha}}$ no longer preserves the polarization, which makes the description of Gérardin’s formula complicated.
What is happening in this example is that the $\theta$-orbit of $\alpha$ is symmetric with respect to the $\Gamma$-action even though $\alpha$ itself is asymmetric. Therefore, to compute each $\operatorname{tr}(\omega_{\ddot{\alpha}}([\eta]^{m_{\alpha}}))$, we have to care about the relationship between the symmetry of the root system $\Phi$ and the twist $\theta$. What we give eventually in this paper is to present an “algorithm” to determine each $\operatorname{tr}(\omega_{\ddot{\alpha}}([\eta]^{m_{\alpha}}))$ based on a case-by-case argument on the type of $\alpha$. Hence, we do not really give a complete description of the twisted character in general. However, depending on the situation ($\mathbf{G}$, $\theta$, and $\mathbf{S}$), the algorithm may terminate rather quickly. For example, it is the case when $\theta$ is involutive.
@@ 2. Notation and assumptions on $p$
Here we summarize our basic notation and assumptions on $p$. (See also the list of symbols at the end of this paper.)
@@ 2.1.1. $p$-adic fields
Let $p$ be a prime number. We fix a $p$-adic field $F$, i.e., $F$ is a finite extension of $\mathbb{Q}_{p}$. We fix an algebraic closure $\overline{F}$ of $F$ and always work within $\overline{F}$ when we consider algebraic extensions of $F$. For any finite extension $E$ of $F$, we write $\mathcal{O}_{E}$, $\mathfrak{p}_{E}$, $k_{E}$, and $\Gamma_{E}$ for the ring of integers of $E$, the maximal ideal of $\mathcal{O}_{E}$, the residue field, and the absolute Galois group $\operatorname{Gal}(\overline{F}/E)$ of $E$, respectively. When $E=F$, we omit the subscript $F$ from these symbols. For any finite extension $E$ of $F$, we write $W_{E}$, $I_{E}$, and $P_{E}$ for the Weil group of $E$, its inertia subgroup, and its wild inertia subgroup, respectively. For any $r\in\mathbb{R}_{>0}$, let $I_{F}^{r}$ denote the $r$-th upper ramification filtration of $I_{F}$. We fix a valuation $\operatorname{val}_{F}$ of $F$ such that $\operatorname{val}_{F}(F^{\times})=\mathbb{Z}$. We extend it to $\overline{F}$ and again write $\operatorname{val}_{F}$ for it. We define an absolute value $|\cdot|_{\overline{F}}$ of $\overline{F}$ by $|\cdot|_{\overline{F}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}q^{-\operatorname{val}_{F}(\cdot)}$, where $q$ is the cardinality of the residue field $k_{F}$.
@@ 2.1.6. Finite sets with Galois actions
We put $\Sigma\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\Gamma\times\{\pm 1\}$. Suppose that $\Phi$ is a finite set with an action of $\Sigma$, e.g., the set of roots of an $F$-rational maximal torus in a connected reductive group (1-1 acts on $\Phi$ via $\alpha\mapsto-\alpha$ in this case). Following [AS09], we put $\dot{\Phi}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\Phi/\Gamma$ and $\ddot{\Phi}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\Phi/\Sigma$. Also, whenever there is no risk of confusion, we simply write $\dot{\alpha}:=\Gamma\alpha$ and $\ddot{\alpha}:=\Sigma\alpha$.
@@
Note that, if $\alpha$ is symmetric, $\dot{\alpha}=\ddot{\alpha}$. This implies that the sets $\dot{\Phi}_{\mathrm{sym}}$ and $\ddot{\Phi}_{\mathrm{sym}}$ can be naturally identified (and, of course, the same is true for $\Phi_{\mathrm{ur}}$ or $\Phi_{\mathrm{ram}}$).
@@ 3.2. Twisted maximal torus
    for any $\eta\in\tilde{\mathbf{S}}$, $\mathbf{G}_{\eta}$ is a connected reductive group with a maximal torus $\mathbf{S}^{\natural}$.
@@ 4.1. Regular supercuspidal representations
The “fibers” of Yu’s construction were investigated by Hakim–Murnaghan; in [HM08], they introduced an equivalence relation called $\mathbf{G}$-equivalence and proved that two cuspidal $\mathbf{G}$-data give rise to isomorphic supercuspidal representations if and only if two data are $\mathbf{G}$-equivalent.
@@ 4.2. Toral supercuspidal representations
Since the depth of $\vartheta$ is $r$, we can extend $\vartheta$ to a character $\hat{\vartheta}$ of $J_{+}$ satisfying $\hat{\vartheta}|_{(S,G)_{\mathbf{x},(r+,s+)}}\equiv\mathbbm{1}$. Then, by the definition of the $\mathbf{G}$-genericity, there exists an element $X^{\ast}$ of $\mathfrak{s}^{\ast}_{-r}$ which is $\mathbf{G}$-generic of depth $r$ in the sense of [Yu01, Section 8] and satisfies
@@ 5.1. Twisted character of a $\theta$-stable representation
$\tilde{\pi}(f)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\int_{\tilde{G}}f(\delta)\tilde{\pi}(\delta)\,d\delta,$